if cos β+cos2β+cos3β=1&sin6β=a+bsin2β=csin4β


find a+b+c


2,3,4,6 are in power..


 


 


 


 


 


 

Grade:12

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

x^3 + x^2 + x = 1

where x is cos (Beta)

x1= 0.54369
x2= -0.77184 + i * 1.11514
x3= -0.77184 - i * 1.11514

Unreal quantities for cosine angles is not possible, Kindly check the question

Regards